OpenStudy (anonymous):

The sum of the first 8 terms of an arithmetic sequence is -20. Find the first 4 terms of the sequence if the 8th term is 8

OpenStudy (anonymous):

an = a1 + (n – 1)d an is the nth term d is the difference between successive terms a1 is the first term the sum is (a1+an)/2

OpenStudy (anonymous):

so for you, an =8 a1 = ? sum =-20 find a1 use it to find d in an = a1 + (n – 1)d

OpenStudy (anonymous):

from there it's easy to find the other terms

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

what did you get for a1?

OpenStudy (anonymous):

is it -4?

OpenStudy (anonymous):

-20 = (8+a1)/2

OpenStudy (anonymous):

multiply both sides by 2 subtract 8 from both sides... a1 =?

OpenStudy (anonymous):

whoops left out an n my bad

OpenStudy (anonymous):

-20=n*(8+a1)/2 so -20 = 8(8+a1)/2 -20/4 = 8+a1 a1=-13

OpenStudy (anonymous):

sorry my algebra skills a little rusty

OpenStudy (anonymous):

mine too evidently. :)

OpenStudy (anonymous):

are you happy with our a1? or do you have questions?

OpenStudy (anonymous):

oh yeah i understand how u got al

OpenStudy (anonymous):

cool.

OpenStudy (anonymous):

show me how to find 'd' now... suing the value for a1 we just found.

OpenStudy (anonymous):

using the value*

OpenStudy (anonymous):

i got 3

OpenStudy (anonymous):

good.

OpenStudy (anonymous):

so first term is -13 second term is -13 +3 third term is -13 +6 and so on...

OpenStudy (anonymous):

so answers are -13, -10, -7 and -4

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

yep:)

OpenStudy (anonymous):

kk cool thanks

OpenStudy (anonymous):

sure!